40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 18 10:46:51 2010, Koji, Update, ABSL, PLL reconstructed 
    Reply  Thu Mar 18 12:27:56 2010, Alberto, Update, ABSL, PLL reconstructed 
       Reply  Thu Mar 18 13:40:35 2010, Koji, Update, ABSL, PLL reconstructed 
          Reply  Thu Mar 18 21:42:26 2010, Koji, Update, ABSL, PLL reconstructed 
             Reply  Mon Mar 22 22:11:26 2010, Mott, Update, ABSL, PLL reconstructed IMG_0150.JPG
                Reply  Mon Mar 22 23:37:32 2010, Mott, Update, ABSL, PLL reconstructed 2010-03-22_23.14.00.jpg2010-03-22_23.24.50.jpg
                   Reply  Tue Mar 23 18:44:46 2010, Mott, Update, ABSL, PLL reconstructed NPRO_PLL_freqresp.png
                      Reply  Wed Mar 24 03:58:18 2010, kiwamu, matt, koji, Update, Green Locking, leave PLL locked DSC_1402.JPG
                         Reply  Wed Mar 24 04:22:51 2010, kiwamu, matt, Update, Green Locking, two NPRO PLL  
                            Reply  Wed Mar 24 14:52:02 2010, Alberto, Update, Green Locking, two NPRO PLL  
                         Reply  Wed Mar 24 15:59:59 2010, kiwamu, matt, Update, Green Locking, leave PLL locked 
                            Reply  Thu Mar 25 17:29:48 2010, kiwamu, mott, Update, Green Locking, PLL two NPROs 
Message ID: 2681     Entry time: Thu Mar 18 13:40:35 2010     In reply to: 2680     Reply to this: 2684
Author: Koji 
Type: Update 
Category: ABSL 
Subject: PLL reconstructed 

We use the current PLL just now, but the renewal of the components are not immediate as it will take some time. Even so we need steady steps towards the better PLL. I appreciate your taking care of it.

Quote:

Quote:

Last night (Mar 17) I checked the PLL setup as Mott had some difficulty to get a clean lock of the PLL setting.

  • I firstly found that the NPRO beam is not going through the Faraday isolator well. This was fixed by aligning the steering mirrors before the Faraday.
     
  • The signal from the RF PD was send to the RF spectrum analyzer through a power splitter. This is a waist of the signal. It was replaced to a directional coupler.
  • Tee-ing the PZT feedback to the oscilloscope was producing the noise in the laser frequency. I put the oscilloscope to the 600Ohm output of the SR560, while connectiong the PZT output to the 50Ohm output.
  • In addition, 6dB+6dB attenuators have been added to the PZT feedback signal.

Now the beating signal is much cleaner and behave straight forward. I will add some numbers such as the PD DC output, RF levels, SR560 settings...

Now I am feeling that we definitely need the development of really clean PLL system as we use PLL everywhere! (i.e. wideband PD, nice electronics, summing amplifiers, stop poking SR560, customize/specialize PDH box, ...etc)

I also had noticed the progressive change of the aux NPRO alignment to the Farady.

I strongly agree about the need of a good and robust PLL.

By modifying the old PDH box (version 2008) eventually I was able to get a PLL robust enough for my purposes. At some point that wasn't good enough for me either.

I then decided to redisign it from scratch. I'm going to work on it. Also because of my other commitments, I'd need a few days/1 week for that. But I'd still like to take care of it. Is it more urgent than that?

 

ELOG V3.1.3-